OpenStudy (anonymous):

I need to create a formula that will tell me how many people are in the US if 1/4th of the people equals 57.7 million. THEN I need to create a formula that will calculate the total number of people in the US next year. HELP?

OpenStudy (anonymous):

if 1/4 = 57.7 million ...lets turn 1/4 into .25 now you need to find the 3/4 of the people. make 3/4 = .75. .25/57.7 = .75/x (cross multiply) .25x = (57.7)(.75) .25x = 43.275 = 43.3 x = 43.3/.25 x = 173.2 million. I am not sure if I did this correctly....If 1/4 is 57.7 then the total population is 173.2 million. However, it is not really a formula, so I am not sure. You might need another opinion.

OpenStudy (anonymous):

Okay Thank you so much!!